You invested a total of $11,200 in shares of the three stocks at the given prices, and expected to earn $304 in annual dividends. If you purchased a total of 250 shares, how many shares of each stock did you purchase?
Price / Dividend Yield
GE(general electric) $16 7%
WMT(Walmart) $56 2%
XOM(Exxon Mobile) $80 2%

Found 2 solutions by MathTherapy, amalm06:
Answer by MathTherapy(10552)


You invested a total of $11,200 in shares of the three stocks at the given prices, and expected to earn $304 in annual dividends. If you purchased a total of 250 shares, how many shares of each stock did you purchase?
Price / Dividend Yield
GE(general electric) $16 7%
WMT(Walmart) $56 2%
XOM(Exxon Mobile) $80 2%

Finance problem on math website!!

Anyway,

Amount of GE shares purchased = 100

Amount of WMT shares purchased = 100

Amount of XOM shares purchased = 50

Answer by amalm06(224)

Number of Shares: x+y+z=250

Cost of Shares: 16x+5y+80z=11200

Dividends: 0.07x+0.02y+0.02z=304

Use Gaussian elimination to solve for x,y, and z.